OpenStudy (rissyroo118):

will medal :} help Which function matches the table? A)ƒ(x) = x + 1 B)ƒ(x) = 1 C)ƒ(x) = x2 D)ƒ(x) = x - 1

OpenStudy (rissyroo118):

X | f(x) 0 | 1 2 | 3 3 | 4 4 | 5 5 | 6 6 | 7

OpenStudy (anonymous):

check which function gives f(x). example x=0 ,f(x)=1 so which function gives 1? then check the next

OpenStudy (rissyroo118):

i think it's D

OpenStudy (rissyroo118):

@Harindu

OpenStudy (anonymous):

lets take a look at D. If you use f(x) = x-1 and plug values in... start with 0. the table tells you that when x=0, f(x) = 1 using equation D, f(0) = 0-1 = -1. They don't match up so it can't be D!

OpenStudy (anonymous):

no your answer is wrong. the function D is f(x)= x-1 so if x=0, f(x) =0-1=-1 which doesn't correspond with the given example

OpenStudy (rissyroo118):

so when i do A it comes out x=1/f-1 b x=1/f c x=^x2/f i believe A comes to the best choice.

OpenStudy (rissyroo118):

@LifeEngineer

OpenStudy (anonymous):

sorry, I'm not following your work. When did you divide by f?

OpenStudy (anonymous):

But you are correct, A is the right answer. Just to recap, B is a constant function. f(x) = 1 is saying that no matter what x is, the function = 1. We know thats not true by looking at every entry but the first in the table. C is saying square your x (unless its 2*x.. either way its wrong) plug in x = 2, x^2 = 4, not 3 so C doesn't work either